# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ A.3 M-Step
The M-step will find the parameters maximizing the function $Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}^{(j)})$ which is defined using $\tau_{i,k}(\boldsymbol{\Theta^{(j)}})$, $\gamma_{i,t,k,h}(\boldsymbol{\Theta^{(j)}})$, $\eta_{i,t,h}(\boldsymbol{\Theta^{(j)}})$ and $\xi_{i,t,k,p,q}(\boldsymbol{\Theta^{(j)}})$.
$Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}^{(j)})$ $=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\boldsymbol{\pi}_{k,\boldsymbol{s}_{i,1}})$
@@
(i) Maximization with respect to $\delta_{k}$:
Consider the parts of (A.3) involving $\delta_{k}$:
$\hat{\delta}^{(j+1)}$ $=$ $\operatorname*{arg\,max}_{\delta}\{\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\delta_{k})\},\qquad s.t.\sum_{k=1}^{K}\delta_{k}=1.$ (18)
Use Lagrange multiplier method to maximize (18) gives us:
@@
$\delta_{k}^{(j+1)}$ $=$ $\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})}{N}.$ (19)
(ii) Update $(\boldsymbol{\Pi}_{k})_{p,q}$:
Consider the parts of (A.3) involving $(\boldsymbol{\Pi}_{k})_{p,q}$:
@@
$s.t.$ $\sum_{q}^{M}(\boldsymbol{\Pi}_{k})_{p,q}=1,\qquad\forall p\in\{1,\dots,M\},k\in\{1,\dots,K\}.$
Similar to the lagrange multiplier method in (19),
$(\boldsymbol{\Pi}_{k})_{p,q}^{(j+1)}=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=2}^{T}\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=2}^{T}\sum_{q^{\prime}=1}^{M}\xi_{i,t,k,p,q^{\prime}}(\boldsymbol{\Theta}^{(j)})}.$ (20)
(iii) Update $(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})_{h=1,\dots,M}$:
Consider the parts in (A.3) involving $(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})_{h=1,\dots,M}$:
@@
Since (21) can be separated into two parts, the $\epsilon_{h}^{(0)},\epsilon_{h}^{(1)}$ and the $a_{h},b_{h}$ can be updated separately.
(iv) Update $(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})_{h=1,\dots,M}$:
$\begin{split}(\hat{\epsilon}_{h}^{0},\hat{\epsilon}_{h}^{1})=&\arg\max_{(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})\\ &[I(x_{i,t}=0)\log(\epsilon_{h}^{(0)})+I(x_{i,t}=1)\log(\epsilon_{h}^{(1)})+I(0<x_{i,t}<1)\log(1-\epsilon_{h}^{(0)}-\epsilon_{h}^{(1)})]\}.\end{split}$ (22)
It is obvious the solutions to (22) are:
@@
(v) Update $(a_{h},b_{h})_{h=1,\dots,M}$:
$(\hat{a}_{h},\hat{b}_{h})=\arg\max_{(a_{h},b_{h})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)\log(f(x_{i,t}\mid a_{h},b_{h}))\}.$ (23)
We utilize the Newton-Raphson method to solve (23). In $r^{th}$ iteration. Use $(a_{h}^{(r)},b_{h}^{(r)})^{\rm T}$ to denote the estimates from last iteration and take derivatives of (23) and denote them as $g_{1}$ and $g_{2}$:
$g_{1}$ $=$ $\frac{\partial\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)\log(f(x_{i,t}\mid a_{h},b_{h}))\}}{\partial a_{h}}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi(a_{h}+b_{h})-\psi(a_{h})+\log(x_{i,t})]\},$
$g_{2}$ $=$ $\frac{\partial\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)\log(f(x_{i,t}\mid a_{h},b_{h}))\}}{\partial b_{h}}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi(a_{h}+b_{h})-\psi(b_{h})+\log(1-x_{i,t})]\}.$
Next, take derivatives of $g_{1}$ and $g_{2}$ , and denote them as $g_{11},g_{12},g_{21},g_{22}$:
$g_{11}$ $=$ $\frac{\partial g_{1}}{\partial a_{h}}=\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}+b_{h})-\psi^{\prime}(a_{h})],$
$g_{12}$ $=$ $\frac{\partial g_{1}}{\partial b_{h}}=\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}+b_{h})],$
$g_{21}$ $=$ $\frac{\partial g_{2}}{\partial a_{h}}=g_{12},$
$g_{22}$ $=$ $\frac{\partial g_{2}}{\partial b_{h}}=\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}+b_{h})-\psi^{\prime}(b_{h})].$
@@
$(a_{h}^{(r+1)},b_{h}^{(r+1)})^{\rm T}$ $=$ $(a_{h}^{(r)},b_{h}^{(r)})^{\rm T}-\boldsymbol{G}^{-1}\boldsymbol{g},$
where $\boldsymbol{G}=\begin{pmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{pmatrix}$, $\boldsymbol{g}=(g_{1},g_{2})^{\rm T}$ and the $\psi$ denotes the Digamma function (Abramowitz and Stegun, 1964), i.e., $\psi(x)=\frac{\partial}{\partial x}\ln{\Gamma(x)}$.
@@ A.4 Pseudo-Code
while $r<\mbox{maxit}_{NR}$ and $\Delta\geq\epsilon_{\mbox{NRtol}}$ do
     Calculate $g_{1},g_{2},g_{11},g_{12},g_{21},g_{22}$ w.r.t. (A.3)
@@
     $\boldsymbol{\Theta}^{(j+1)}$ is the collection of all new parameters from M-step.
until Convergence of log-likelihood or max iteration reached.
Output: Estimated hidden states, cluster information, and $\widehat{\boldsymbol{\Theta}}$, the MLE of parameters.
A.5 Variance Estimation
@@
Input: The MLE of parameters, $\widehat{\boldsymbol{\Theta}}$, and the estimated parameters in $j$-th iteration, $\boldsymbol{\Theta}^{(j)}$.
for $p=1$ to $|\widehat{\boldsymbol{\Theta}}|$ do
     Write $\widehat{\boldsymbol{\Theta}}^{(j)}_{p}$, the one-step-rollback estimation on the $p$-th parameter in the $\widehat{\boldsymbol{\Theta}}$.
     Treat $\widehat{\boldsymbol{\Theta}}^{(j)}_{p}$ as current estimation, and run one iteration of EM to obtain $\widetilde{\boldsymbol{\Theta}}^{(j+1)}_{p}$.
@@
As the second key part in (24), $\boldsymbol{I}^{-1}_{oc}$ is the inverse of the observed data information matrix evaluated at MLE. Similar to the situation discussed in Meng and Rubin (1991), the direct computation of the observed data information matrix is very difficult. We thus first calculate the complete data information matrix $\boldsymbol{I}_{c}$ and then take its expectation over the conditional distribution $f(\boldsymbol{S},\boldsymbol{Z}\mid\boldsymbol{X},\boldsymbol{\Theta})$ evaluated at $\boldsymbol{\Theta}=\widehat{\boldsymbol{\Theta}}$. The complete data information matrix $\boldsymbol{I}_{c}$ is calculated by $\boldsymbol{I}_{c}(\boldsymbol{\Theta}\mid\boldsymbol{X,S,Z})=-\frac{\partial^{2}\log f(\boldsymbol{X,S,Z}\mid\boldsymbol{\Theta})}{\partial\boldsymbol{\Theta}^{2}}$, which fully depends on the complete data likelihood. We have that $\boldsymbol{I}_{oc}=E[\boldsymbol{I}_{c}(\boldsymbol{\Theta}\mid\boldsymbol{X},\boldsymbol{S},\boldsymbol{Z})\mid\boldsymbol{X},\boldsymbol{\Theta}]|_{\boldsymbol{\Theta}=\widehat{\boldsymbol{\Theta}}}$. That is, $\boldsymbol{\Theta}=(\epsilon_{1}^{(0)},\epsilon_{1}^{(1)},a_{1},b_{1},\ldots,\epsilon_{M}^{(0)},\epsilon_{M}^{(1)},a_{M},b_{M})$, and $\boldsymbol{I}_{oc}$ is in a format of sparse block diagonal matrix,
$\begin{Bmatrix}\boldsymbol{I}_{oc}^{(1)}&0&0&0\\ 0&\boldsymbol{I}_{oc}^{(2)}&0&0\\ 0&0&\cdots&0\\ 0&0&0&\boldsymbol{I}_{oc}^{(M)}\end{Bmatrix}$

### v2

@@ A.3 M-Step
In the M-step, the parameters are updated by maximizing the expected complete-data log-likelihood $Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}^{(j)})$, which has been reformulated in terms of the intermediate quantities $\tau_{i,k}(\boldsymbol{\Theta}^{(j)})$, $\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})$, $\eta_{i,t,h}(\boldsymbol{\Theta}^{(j)})$ and $\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})$.
$Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}^{(j)})$ $=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\delta_{k})+\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\boldsymbol{\pi}_{k,\boldsymbol{s}_{i,1}})$
@@
(i) Maximize with respect to $\delta_{k}$:
$\hat{\boldsymbol{\delta}}^{(j+1)}$ $=$ $\operatorname*{arg\,max}_{\boldsymbol{\delta}}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\delta_{k}),\qquad s.t.\sum_{k=1}^{K}\delta_{k}=1.$ (11)
Applying the method of Lagrange multipliers to maximize (11) leads to the following derivation:
@@
$\delta_{k}^{(j+1)}$ $=$ $\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})}{N}.$
(ii) Maximization with respect to $\boldsymbol{\pi}_{k,\boldsymbol{s}_{i,1}}$:
$\widehat{\boldsymbol{\pi}}^{(j+1)}$ $=$ $\operatorname*{arg\,max}_{\boldsymbol{\pi}}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\log(\hat{\boldsymbol{\pi}}^{(j)}),$
$s.t.$ $\sum_{h=1}^{M}\boldsymbol{\pi}_{k,h}=1,\forall k\in\{1,\dots,K\}.$
Using the method of Lagrange multipliers, the result is
$\widehat{\boldsymbol{\pi}}_{k,h}^{(j+1)}=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\gamma_{i,1,k,h}(\boldsymbol{\Theta}^{(j)})}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})}.$
(iii) Maximization with respect to $(\boldsymbol{\Pi}_{k})_{p,q}$:
@@
$s.t.$ $\sum_{q=1}^{M}(\boldsymbol{\Pi}_{k})_{p,q}=1,\qquad\forall p\in\{1,\dots,M\},k\in\{1,\dots,K\}.$
Using the method of Lagrange multipliers, the result is
$(\boldsymbol{\Pi}_{k})_{p,q}^{(j+1)}=\frac{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=2}^{T}\xi_{i,t,k,p,q}(\boldsymbol{\Theta}^{(j)})}{\sum_{i=1}^{N}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\sum_{t=2}^{T}\sum_{q^{\prime}=1}^{M}\xi_{i,t,k,p,q^{\prime}}(\boldsymbol{\Theta}^{(j)})}.$
(iv) Maximization with respect to $(a_{h},b_{h},\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})_{h=1,\dots,M}$:
@@
The maximization problem in (12) can be decomposed into two components, $\epsilon_{h}^{(0)},\epsilon_{h}^{(1)}$ and the $a_{h},b_{h}$, which can be updated separately.
(v) Update $(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})_{h=1,\dots,M}$:
$\begin{split}(\hat{\epsilon}_{h}^{(0)},\hat{\epsilon}_{h}^{(1)}=&\operatorname*{arg\,max}_{(\epsilon_{h}^{(0)},\epsilon_{h}^{(1)})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})\\ &[I(x_{i,t}=0)\log(\epsilon_{h}^{(0)})+I(x_{i,t}=1)\log(\epsilon_{h}^{(1)})+I(0<x_{i,t}<1)\log(1-\epsilon_{h}^{(0)}-\epsilon_{h}^{(1)})]\}.\end{split}$ (13)
The solutions to (13) are:
@@
(vi) Update $(a_{h},b_{h})_{h=1,\dots,M}$:
$(\hat{a}_{h},\hat{b}_{h})=\operatorname*{arg\,max}_{(a_{h},b_{h})}\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)\log(f(x_{i,t}\mid a_{h},b_{h}))\}.$ (14)
We employ the Newton-Raphson method to solve (14). Let $(a_{h}^{(r)},b_{h}^{(r)})^{\rm T}$ denote the parameter estimates obtained at $r^{\mbox{th}}$ iteration. The gradient vector is defined as $\boldsymbol{g}=(g_{1},g_{2})^{\rm T}$, where $g_{1}$ and $g_{2}$ denote the first-order partial derivatives of the objective function in (14) with respect to $a_{h}^{(r)}$ and $b_{h}^{(r)}$, respectively. The Hessian matrix $\boldsymbol{G}$ is defined as $\boldsymbol{G}=\begin{pmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{pmatrix}$, where each $g_{ij}$ denotes the second-order partial derivative with respect to the corresponding parameters. To simplify the notation in the subsequent derivations, we denote the objective function in (14) by $\mathcal{L}$ and let the $\psi$ denote the Digamma function, i.e., $\psi(x)=\frac{\partial}{\partial x}\ln{\Gamma(x)}$.
$g_{1}$ $=$ $\frac{\partial\mathcal{L}}{\partial a_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi(a_{h}^{(r)}+b_{h}^{(r)})-\psi(a_{h}^{(r)})+\log(x_{i,t})]\},$
$g_{2}$ $=$ $\frac{\partial\mathcal{L}}{\partial b_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi(a_{h}^{(r)}+b_{h}^{(r)})-\psi(b_{h}^{(r)})+\log(1-x_{i,t})]\},$
$g_{11}$ $=$ $\frac{\partial g_{1}}{\partial a_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}^{(r)}+b_{h}^{(r)})-\psi^{\prime}(a_{h}^{(r)})],$
$g_{12}$ $=$ $\frac{\partial g_{1}}{\partial b_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}^{(r)}+b_{h}^{(r)})],$
$g_{21}$ $=$ $\frac{\partial g_{2}}{\partial a_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}=g_{12},$
$g_{22}$ $=$ $\frac{\partial g_{2}}{\partial b_{h}}|_{(a_{h},b_{h})=(a_{h}^{(r)},b_{h}^{(r)})}$
$=$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\tau_{i,k}(\boldsymbol{\Theta}^{(j)})\{\sum_{t=1}^{T}\gamma_{i,t,k,h}(\boldsymbol{\Theta}^{(j)})I(0<x_{i,t}<1)[\psi^{\prime}(a_{h}^{(r)}+b_{h}^{(r)})-\psi^{\prime}(b_{h}^{(r)})].$
@@
$(a_{h}^{(r+1)},b_{h}^{(r+1)})^{\rm T}$ $=$ $(a_{h}^{(r)},b_{h}^{(r)})^{\rm T}-\boldsymbol{G}^{-1}\boldsymbol{g}.$
@@ A.4 Pseudo-Code
while $r<\mbox{maxit}_{\mbox{NR}}$ and $\Delta\geq\epsilon_{\mbox{NRtol}}$ do
     Calculate $g_{1},g_{2},g_{11},g_{12},g_{21},g_{22}$
@@
     $\boldsymbol{\Theta}^{(j+1)}$ is the collection of all new parameters from M-step
until Convergence of log-likelihood or max iteration reached
Output: Estimated hidden states, estimated cluster and $\widehat{\boldsymbol{\Theta}}$, the MLE of parameters.
Appendix B Inference
B.1 Variance Estimation
@@
Input: The MLE of parameters, $\widehat{\boldsymbol{\Phi}}$, and the estimated parameters in $j$-th iteration, $\boldsymbol{\Phi}^{(j)}$.
for $p=1$ to $|\widehat{\boldsymbol{\Phi}}|$ do
     Write $\widehat{\boldsymbol{\Phi}}^{(j)}_{p}$, the one-step-rollback estimation on the $p$-th parameter in the $\widehat{\boldsymbol{\Phi}}$.
     Treat $\widehat{\boldsymbol{\Phi}}^{(j)}_{p}$ as current estimation, and run one iteration of EM to obtain $\widetilde{\boldsymbol{\Phi}}^{(j+1)}_{p}$.
@@
As the second key part in (15), $\boldsymbol{I}^{-1}_{oc}$ is the inverse of the observed data information matrix evaluated at MLE. Similar to the situation discussed in Meng and Rubin (1991), considering the direct computation of the observed-data information matrix is very difficult, we firstly calculate the complete-data information matrix $\boldsymbol{I}_{c}$ and then take its expectation over the conditional distribution $f(\boldsymbol{S},\boldsymbol{Z}\mid\boldsymbol{X},\boldsymbol{\Phi})$ evaluated at $\boldsymbol{\Phi}=\widehat{\boldsymbol{\Phi}}$. The complete data information matrix $\boldsymbol{I}_{c}$ is calculated by $\boldsymbol{I}_{c}(\boldsymbol{\Theta}\mid\boldsymbol{X,S,Z})=-\frac{\partial^{2}\log f(\boldsymbol{X,S,Z}\mid\boldsymbol{\Theta})}{\partial\boldsymbol{\Theta}^{2}}$, which fully depends on the complete data likelihood. And $\boldsymbol{I}_{oc}=E[\boldsymbol{I}_{c}(\boldsymbol{\Theta}\mid\boldsymbol{X},\boldsymbol{S},\boldsymbol{Z})\mid\boldsymbol{X},\boldsymbol{\Phi}]|_{\boldsymbol{\Phi}=\widehat{\boldsymbol{\Phi}}}$. That is, in our case, $\boldsymbol{\Phi}=(\epsilon_{1}^{(0)},\epsilon_{1}^{(1)},a_{1},b_{1},\ldots,\epsilon_{M}^{(0)},\epsilon_{M}^{(1)},a_{M},b_{M})$ and $\boldsymbol{I}_{oc}$ can be represented as a sparse matrix,
$\begin{bmatrix}\boldsymbol{I}_{oc}^{(1)}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{I}_{oc}^{(2)}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\cdots&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}&\boldsymbol{I}_{oc}^{(M)}\end{bmatrix}$